### v2

# Diophantine Approximation with Piatetski-Shapiro Primes

Abstract.
We prove that for every irrational number $\alpha$, real number $\beta$, real number $c$ satisfying $81<c<9/$ 8 and positive real number $\theta$ satisfying $\theta<(9/c-8)/10$, there exist infinitely many primes of the form $p=\left[n^{c}\right]$ with $n\in\mathbb{N}$ such that $||\alpha p+\beta||<p^{-\theta}$.
Key words and phrases:
Piatetski-Shapiro primes, rational approximation with prime denominator, Harman’s sieve. exponential sums
2010 Mathematics Subject Classification:
11J25,11J54,11J71,11L07,11L20,11N36
1. Introduction
In this article, we consider Diophantine approximation with denominators which are restricted to Piatetski-Shapiro primes. These are prime numbers $p$ of the form $\left[n^{c}\right]$, where $c>1$ is a fixed real number, $n$ runs over the positive integers, and $[x]$ denotes the integral part of $x\in\mathbb{R}$. One may conjecture that, given any non-integer $c>1$, there exist infinitely many such primes. Their investigation was initiated by Piatetski-Shapiro [9] who demonstrated their infinitude for $1.09091<c<12/11=$ 1.0909 …. This range has been widened by many authors. The latest record is due to Rivat and Wu [10] who obtained a range of $1.18531<c<243/205=$ 1.1853 …. Progress on Piatetski-Shapiro primes measures progress on exponential sums and sieve methods.
The Dirichlet approximation theorem, a cornerstone in Diophantine approximation, implies that for any irrational number $\alpha$, there exist infinitely many positive integers $q$ such that $||\alpha q||<q^{-1}$, where $||x||$ is the distance of $x\in\mathbb{R}$ to the nearest integer. Interesting problems arise when the $q$’s are restricted to a sparse subset of the natural numbers. Many authors have considered the question for which $\theta>0$ one can show the infinitude of primes $p$ such that $||\alpha p||<p^{-\theta}$. It may be expected that an exponent of $\theta=1-\varepsilon$ is admissible. The first author to consider this problem was Vinogradov who proved that $\theta=1/5-\varepsilon$ is admissible [11]. The current record is due to Matomäki [8] who obtained an exponent of $\theta=1/3-\varepsilon$. Harman’s sieve has become a standard tool for the investigation of this problem (for details on Harman’s sieve see [5]).
Dimitrov considered a hybrid problem, restricting the set of primes $p$ in the above Diophantine approximation problem to Piatetski-Shapiro primes. He proved the following in [2].
Theorem 1.1 (Dimitrov).
Fix an irrational number $\alpha$, a real number $\beta$, a real number $c$ satisfying $111<c<12/$ 11 and a positive real number $\theta$ satisfying $\theta<(12/c-11)/26$. Then there exist infinitely many primes of the form $p=\left[n^{c}\right]$ with $n\in\mathbb{N}$ such that $||\alpha p+\beta||<p^{-\theta}$.
In this article, we improve Dimitrov’s result, widening the $c$- and $\theta$-ranges. We will establish the following.
Theorem 1.2.
Fix an irrational number $\alpha$, a real number $\beta$, a real number $c$ satisfying $81<c<9/$ 8 and a positive real number $\theta$ satisfying $\theta<(9/c-8)/10$. Then there exist infinitely many primes of the form $p=\left[n^{c}\right]$ with $n\in\mathbb{N}$ such that $||\alpha p+\beta||<p^{-\theta}$.
Whereas Dimitrov used Vaughan’s identity to obtain his result, we here apply Harman’s sieve. We also employ arguments due to Heath-Brown [7] and Balog and Friedlander [1] to improve certain estimates of bilinear exponential sums.
Acknowledgements. The first-named author would like to thank the Ramakrishna Mission Vivekananda Educational and Research Insititute for an excellent work environment. The research of the second-named author was supported by a Prime Minister Research Fellowship (PMRF ID- 0501972), funded by the Ministry of Education, Govt. of India.
2. Simplification and application of Harman’s sieve
Let us denote by $\mathbb{P}$ the set of prime numbers. Our goal is to prove that
$\sharp\left\{p\in\mathbb{P}:X/2\leq p<X,\ p=\left[n^{c}\right]\mbox{ for some }n\in\mathbb{N},\ ||\alpha p+\beta||<\Delta\right\}>0$
for infinitely many $X\in\mathbb{N}$, where we set
$\Delta:=X^{-\theta}.$ (2.1)
Let $\gamma=1/c$. Obviously, $p=\left[n^{c}\right]$ for some $n\in\mathbb{N}$ if and only if
$p^{\gamma}\leq n<(p+1)^{\gamma}\quad\mbox{for some }n\in\mathbb{N}.$ (2.2)
Moreover, $(p+1)^{\gamma}=p^{\gamma}+\gamma p^{\gamma-1}+O\left(p^{\gamma-2}\right)$. Hence, if
$\delta:=\frac{\gamma X^{\gamma-1}}{10},$ (2.3)
then
$p^{\gamma}\in(n-3\delta,n-\delta)\quad\mbox{for some }n\in\mathbb{N}$ (2.4)
implies (2.2), provided that $X/2\leq p<X$ with $X$ large enough. Clearly, the condition (2.4) is equivalent to the inequality
$||p^{\gamma}+2\delta||<\delta.$ (2.5)
Therefore, it suffices to show that for infinitely many $X\in\mathbb{N}$, we have
$\sharp(\mathcal{A}\cap\mathbb{P})>0,$ (2.6)
where we set
$\mathcal{A}:=\left\{a\in\mathbb{N}:X/2\leq a<X,\ ||\alpha a+\beta||<\Delta,\ ||a^{\gamma}+2\delta||<\delta\right\}.$ (2.7)
By the Dirichlet approximation theorem, mentioned in section 1, there exist infinitely many positive integers $q$ such that
$\left|\alpha-\frac{a}{q}\right|<q^{-2}\quad\mbox{for some }a\in\mathbb{Z}\mbox{ with }(a,q)=1.$ (2.8)
We shall prove that (2.6) holds if $X$ is a positive integer such that
$X^{2\theta+10\eta}\leq q\leq X^{1-\theta-10\eta}$ (2.9)
for some $q\in\mathbb{N}$ satisfying (2.8), where $\eta$ is a suitably small positive number. Since there are infinitely many such positive integers $X$, the result in Theorem 1.2 then follows.
To establish (2.6), we shall employ Harman’s sieve. Set
$\mathcal{B}:=[X/2,X)\cap\mathbb{N}.$ (2.10)
The idea is to compare the cardinalities $\sharp(\mathcal{A}\cap\mathbb{P})$ and $\sharp(\mathcal{B}\cap\mathbb{P})$ via a comparison of bilinear sums associated to the sets $\mathcal{A}$ and $\mathcal{B}$. Precisely, we have the following.
Proposition 2.1 (Harman).
Let $X\geq 4$ and $\mathcal{A}\subseteq\mathcal{B}:=[X/2,X)\cap\mathbb{N}$. Suppose that for some $\eta>0$ and some $\lambda>0$ we have, for all sequences $a_{m}$ and $b_{n}$ of non-negative real numbers satisfying
$a_{m}\leq\tau(m),\ b_{n}\leq\tau(n)$ (2.11)
($\tau(n)$ here denotes the number of divisors of $n$) that
$\sum_{\begin{subarray}{c}mn\in\mathcal{A}\\ m\leq X^{15/22}\end{subarray}}a_{m}=\lambda\sum_{\begin{subarray}{c}mn\in\mathcal{B}\\ m\leq X^{15/22}\end{subarray}}a_{m}+O\left(\lambda X^{1-\eta}\right)$ (2.12)
and
$\sum_{\begin{subarray}{c}mn\in\mathcal{A}\\ X^{7/22}\leq m\leq X^{8/22}\end{subarray}}a_{m}b_{n}=\lambda\sum_{\begin{subarray}{c}mn\in\mathcal{B}\\ X^{7/22}\leq m\leq X^{8/22}\end{subarray}}a_{m}b_{n}+O\left(\lambda X^{1-\eta}\right).$ (2.13)
Then the inequality
$\sharp(\mathcal{A}\cap\mathbb{P})>\frac{\lambda}{10}\cdot\sharp(\mathcal{B}\cap\mathbb{P})$
holds, provided that $X$ is large enough.
Proof.
Apply [4, Theorem 2] with $\theta=7/22$ and note that if $X\geq 4$, then $\sharp(\mathcal{A}\cap\mathbb{P})=S(\mathcal{A},X^{1/2})$ and $\sharp(\mathcal{B}\cap\mathbb{P})=S(\mathcal{B},X^{1/2})$, where the sieve function $S(\mathcal{C},z)$ is defined as in [4]. Also note that the sequences $a_{m}$ and $b_{n}$ can be confined to non-negative real numbers. (In the original statement, $a_{m}$ and $b_{n}$ were assumed to be complex. Extracting the real and imaginary parts of $a_{m}$ and $b_{n}$, it suffices to assume that they are real. Then dividing the sums over $m$ and $n$ into subsums according to the signs of $a_{m}$ and $b_{n}$, it suffices to assume that they are non-negative.) ∎
Following usual terminology, we call the sums in (2.12) type I and the sums in (2.13) type II sums.
3. Fourier analysis
We shall detect the conditions $||\alpha a+\beta||<\Delta$ and $||a^{\gamma}+2\delta||<\delta$ in (2.7) using the following Fourier analytic device.
Lemma 3.1 (Harman).
Let $\xi\in(0,1)$ and $K$ be any positive integer. Define
$\chi_{\xi}(x)=\begin{cases}1,\ \text{ if }||x||<\xi,\\ 0,\ \text{otherwise.}\end{cases}$
Then there are sequences $c^{-}_{k}$ and $c^{+}_{k}$ of complex numbers such that
$|c^{\pm}_{k}|\leq\min\left\{2\xi+\frac{1}{K+1},\ \frac{3}{2|k|}\right\}$
for all $k\in\mathbb{N}$ and
$\chi_{\xi}^{-}(x)\leq\chi_{\xi}(x)\leq\chi_{\xi}^{+}(x)$
for all $x\in\mathbb{R}$, where
$\chi_{\xi}^{-}(x):=2\xi-\frac{1}{K+1}-\sum_{0<|k|\leq K}c^{-}_{k}e(kx)\quad\mbox{and}\quad\chi_{\xi}^{+}(x):=2\xi+\frac{1}{K+1}+\sum_{0<|k|\leq K}c^{+}_{k}e(kx).$
(Here we define $e(z):=e^{2\pi iz}$ for all $z\in\mathbb{R}$).
Proof.
This is [5, Lemma 2.1]. ∎
We recall the definitions of $\Delta$ and $\delta$ in (2.1) and (2.3), fix an arbitrarily small $\eta>0$ and set
$L:=\left[\Delta^{-1}X^{\eta}\right]=\left[X^{\theta+\eta}\right]\quad\mbox{and}\quad H:=\left[\delta^{-1}X^{\eta}\right]=\left[10X^{1-\gamma+\eta}/\gamma\right].$ (3.1)
Then Lemma 3.1 above with $\xi=\Delta,\delta$ and $K=L,H$ produces functions $\chi_{\Delta}^{\pm}$ and $\chi_{\delta}^{\pm}$ such that
$\chi_{\Delta}^{-}(x)\leq\chi_{\Delta}(x)\leq\chi_{\Delta}^{+}(x)$
and
$\chi_{\delta}^{-}(y)\leq\chi_{\delta}(y)\leq\chi_{\delta}^{+}(y)$
for all $x,y\in\mathbb{R}$. We want to use these functions to bound the product $\chi_{\Delta}(x)\chi_{\delta}(y)$ from below and above. To this end, we use the following observation from [6]: For all $x,y\in\mathbb{R}$, we have
$\Xi^{-}(x,y)\leq\chi_{\Delta}(x)\chi_{\delta}(y)\leq\Xi^{+}(x,y),$
where
$\Xi^{-}(x,y):=\chi_{\Delta}^{-}(x)\chi_{\delta}^{+}(y)+\chi_{\Delta}^{+}(x)\chi_{\delta}^{-}(y)-\chi_{\Delta}^{+}(x)\chi_{\delta}^{+}(y)\quad\mbox{and}\quad\Xi^{+}(x,y):=\chi_{\Delta}^{+}(x)\chi_{\delta}^{+}(y).$
Recalling the definitions of the sets $\mathcal{A}$ and $\mathcal{B}$ in (2.7) and (2.10), and keeping the condition of non-negativity of $a_{m}$ and $b_{n}$ in Proposition 2.1 in mind, it follows that the sums on the left-hand sides of (2.12) and (2.13) are bounded from below and above by
$\sum_{\begin{subarray}{c}mn\in\mathcal{B}\\ m\leq X^{15/22}\end{subarray}}a_{m}\Xi^{-}\left(\alpha mn+\beta,(mn)^{\gamma}+2\delta\right)\leq\sum_{\begin{subarray}{c}mn\in\mathcal{A}\\ m\leq X^{15/22}\end{subarray}}a_{m}\leq\sum_{\begin{subarray}{c}mn\in\mathcal{B}\\ m\leq X^{15/22}\end{subarray}}a_{m}\Xi^{+}\left(\alpha mn+\beta,(mn)^{\gamma}+2\delta\right)$
and
$\begin{split}\sum_{\begin{subarray}{c}mn\in\mathcal{B}\\ X^{7/22}\leq m\leq X^{8/22}\end{subarray}}a_{m}b_{n}\Xi^{-}\left(\alpha mn+\beta,(mn)^{\gamma}+2\delta\right)\leq&\sum_{\begin{subarray}{c}mn\in\mathcal{A}\\ X^{7/22}\leq m\leq X^{8/22}\end{subarray}}a_{m}b_{n}\\ \leq&\sum_{\begin{subarray}{c}mn\in\mathcal{B}\\ X^{7/22}\leq m\leq X^{8/22}\end{subarray}}a_{m}b_{n}\Xi^{+}\left(\alpha mn+\beta,(mn)^{\gamma}+2\delta\right).\end{split}$
Now using the definitions of $\Xi^{-}(x,y)$ and $\Xi^{+}(x,y)$, replacing $\chi_{\Delta}^{\pm}(x)$ and $\chi_{\delta}^{\pm}(y)$ by the relevant trigonometrical polynomials from Lemma 3.1, multiplying out, and noting the condition (2.11) and the well-known bound $\tau(n)\ll_{\varepsilon}n^{\varepsilon}$ for every $\varepsilon>0$, we deduce that
$\sum_{\begin{subarray}{c}mn\in\mathcal{A}\\ m\leq X^{15/22}\end{subarray}}a_{m}=\lambda\sum_{\begin{subarray}{c}mn\in\mathcal{B}\\ m\leq X^{15/22}\end{subarray}}a_{m}+O\left(\lambda X^{1-\eta}+\Sigma_{I}\right)$
and
$\sum_{\begin{subarray}{c}mn\in\mathcal{A}\\ X^{7/22}\leq m\leq X^{8/22}\end{subarray}}a_{m}b_{n}=\lambda\sum_{\begin{subarray}{c}mn\in\mathcal{B}\\ X^{7/22}\leq m\leq X^{8/22}\end{subarray}}a_{m}b_{n}+O\left(\lambda X^{1-\eta}+\Sigma_{II}\right)$
with
$\lambda:=4\Delta\delta,$ (3.2)
where the error term $\Sigma_{1}$ is a sum of expressions of the form
$\mathcal{S}_{1}:=\delta\sum_{\begin{subarray}{c}mn\sim X\\ m\leq X^{15/22}\end{subarray}}\sum_{0<|l|\leq L}a_{m}c_{l}e(\alpha lmn),$ (3.3)
$\mathcal{S}_{2}:=\Delta\sum_{\begin{subarray}{c}mn\sim X\\ m\leq X^{15/22}\end{subarray}}\sum_{0<|h|\leq H}a_{m}d_{h}e\left(h(mn)^{\gamma}\right),$ (3.4)
$\mathcal{S}_{3}:=\sum_{\begin{subarray}{c}mn\sim X\\ m\leq X^{15/22}\end{subarray}}\sum_{0<|l|\leq L}\sum_{0<|h|\leq H}a_{m}c_{l}d_{h}e\left(\alpha lmn+h(mn)^{\gamma}\right),$ (3.5)
and the error term $\Sigma_{2}$ is a sum of expressions of the form
$\mathcal{T}_{1}:=\delta\sum_{\begin{subarray}{c}mn\sim X\\ X^{7/22}\leq m\leq X^{8/22}\end{subarray}}\sum_{0<|l|\leq L}a_{m}b_{n}c_{l}e(\alpha lmn),$ (3.6)
$\mathcal{T}_{2}:=\Delta\sum_{\begin{subarray}{c}mn\sim X\\ X^{7/22}\leq m\leq X^{8/22}\end{subarray}}\sum_{0<|h|\leq H}a_{m}b_{n}d_{h}e\left(h(mn)^{\gamma}\right),$ (3.7)
$\mathcal{T}_{3}:=\sum_{\begin{subarray}{c}mn\sim X\\ X^{7/22}\leq m\leq X^{8/22}\end{subarray}}\sum_{0<|l|\leq L}\sum_{0<|h|\leq H}a_{m}b_{n}c_{l}d_{h}e\left(\alpha lmn+h(mn)^{\gamma}\right),$ (3.8)
the coefficients satisfying the bounds
$a_{m}\ll m^{\varepsilon},\quad b_{n}\ll n^{\varepsilon},\quad c_{l}\ll\Delta,\quad d_{h}\ll\delta.$
Here and in the following, the notation $x\sim X$ means that $X/2\leq x<X$. In view of Proposition 2.1 and the prime number theorem, to establish Theorem 1.2, it now suffices to prove that
$\mathcal{S}_{i},\mathcal{T}_{i}\ll\lambda X^{1-\eta}$
for $i=1,2,3$ and a suitably small $\eta>0$. This is the content of the remainder of this paper.
4. Tailoring the type I and II sums
It will be advantageous to tailor the expressions $\mathcal{S}_{i}$, $\mathcal{T}_{i}$ above. Breaking the sums on the right-hand sides of (3.3)-(3.8) into dyadic subsums with $m\sim M$, $n\sim N$, $|l|\sim U$, $|h|\sim V$ and scaling the coefficients, setting
$a_{m}^{\ast}:=\frac{a_{m}}{m^{\varepsilon}},\quad b_{n}^{\ast}:=\frac{b_{n}}{n^{\varepsilon}},\quad c_{l}^{\ast}:=\frac{c_{l}}{\Delta},\quad d_{h}^{\ast}:=\frac{d_{h}}{\delta},$
we reduce these sums to $O(\log^{4}X)$ expressions of the form
$\mathcal{S}_{1}^{\ast}:=\sum_{\begin{subarray}{c}m\sim M\\ n\sim N\\ mn\sim X\end{subarray}}\sum_{|l|\sim U}a_{m}^{\ast}c_{l}^{\ast}e(\alpha lmn),$ (4.1)
$\mathcal{S}_{2}^{\ast}:=\sum_{\begin{subarray}{c}m\sim M\\ n\sim N\\ mn\sim X\end{subarray}}\sum_{|h|\sim V}a_{m}^{\ast}d_{h}^{\ast}e\left(h(mn)^{\gamma}\right),$ (4.2)
$\mathcal{S}_{3}^{\ast}:=\sum_{\begin{subarray}{c}m\sim M\\ n\sim N\\ mn\sim X\end{subarray}}\sum_{|l|\sim U}\sum_{|h|\sim V}a_{m}^{\ast}c_{l}^{\ast}d_{h}^{\ast}e\left(\alpha lmn+h(mn)^{\gamma}\right),$ (4.3)
$\mathcal{T}_{1}^{\ast}:=\sum_{\begin{subarray}{c}m\sim M\\ n\sim N\\ mn\sim X\end{subarray}}\sum_{|l|\sim U}a_{m}^{\ast}b_{n}^{\ast}c_{l}^{\ast}e(\alpha lmn),$ (4.4)
$\mathcal{T}_{2}^{\ast}:=\sum_{\begin{subarray}{c}m\sim M\\ n\sim N\\ mn\sim X\end{subarray}}\sum_{|h|\sim V}a_{m}^{\ast}b_{n}^{\ast}d_{h}^{\ast}e\left(h(mn)^{\gamma}\right),$ (4.5)
$\mathcal{T}_{3}^{\ast}:=\sum_{\begin{subarray}{c}m\sim M\\ n\sim N\\ mn\sim X\end{subarray}}\sum_{|l|\sim U}\sum_{|h|\sim V}a_{m}^{\ast}b_{n}^{\ast}c_{l}^{\ast}d_{h}^{\ast}e\left(\alpha lmn+h(mn)^{\gamma}\right),$ (4.6)
where the coefficients $a_{m}^{\ast}$, $b_{n}^{\ast}$, $c_{l}^{\ast}$, $d_{h}^{\ast}$ are complex numbers satisfying
$a_{m}^{\ast},b_{n}^{\ast},c_{l}^{\ast},d_{h}^{\ast}\ll 1.$
Now it remains to prove that
$\mathcal{S}_{i}^{\ast}\ll X^{1-2\eta}\mbox{ for }i=1,2,3,\ M\leq X^{15/22},\ U\leq L,\ V\leq H$ (4.7)
and
$\mathcal{T}_{i}^{\ast}\ll X^{1-2\eta}\mbox{ for }i=1,2,3,\ X^{7/22}\leq M\leq X^{8/22},\ N\asymp X/M,\ U\leq L,V\leq H$ (4.8)
for $\eta>0$ small enough. Here and throughout the sequel, for $x,y>0$, the notation $x\asymp y$ means that there are constants $C_{2}>C_{1}>0$ such that $C_{1}x\leq y\leq C_{2}x$.
5. Estimations of $\mathcal{S}_{1}^{\ast}$ and $\mathcal{T}_{1}^{\ast}$
Recall (2.8) and (2.9). To estimate $\mathcal{S}_{1}^{\ast}$ and $\mathcal{T}_{1}^{\ast}$, defined in (4.1) and (4.4), we use the following standard bounds for sums involving linear exponential terms.
Lemma 5.1.
Let $K,N\geq 1$ and $\alpha_{k},\beta_{n}$ be any sequences of complex numbers. Assume that (2.8) holds. Then we have
$\sum\limits_{n\asymp N}e(\alpha kn)\ll\min\left\{N,||\alpha k||^{-1}\right\}\mbox{ for all }k\in\mathbb{N},$ (5.1)
$\sum\limits_{k\asymp K}\min\left\{N,||\alpha k||^{-1}\right\}\ll\left(\frac{KN}{q}+K+q\right)(\log 2KNq)$ (5.2)
and
$\sum_{\begin{subarray}{c}k\asymp K\\ n\asymp N\\ kn\asymp KN\end{subarray}}\alpha_{k}\beta_{n}e(\alpha mn)\ll\left(\sum_{k\asymp K}|\alpha_{k}|^{2}\sum_{n\asymp N}|\beta_{n}|^{2}\right)^{1/2}\left(\frac{KN}{q}+K+N+q\right)^{1/2}(\log 2KNq)^{1/2}.$ (5.3)
Proof.
See [5, section 1.6]. ∎
Now recalling the inequalities
$\frac{8}{9}<\gamma=1/c<1\quad\mbox{and}\quad 0<\theta<\frac{9\gamma-8}{10}<\frac{1}{10}$ (5.4)
from Theorem 1.2, we are ready to prove the desired bounds for $\mathcal{S}_{1}^{\ast}$ and $\mathcal{T}_{1}^{\ast}$. Throughout the sequel, we assume that $\varepsilon$ is a fixed but arbitrarily small positive real number.
Lemma 5.2.
There is $\eta>0$ such that $\mathcal{S}_{1}^{\ast},\mathcal{T}_{1}^{\ast}\ll X^{1-2\eta}$.
Proof.
Applying (5.1) to the sum over $n$ on the right-hand side of (4.1), we have
$\mathcal{S}_{1}^{\ast}\ll\sum_{m\sim M}\sum_{l\sim U}a_{m}^{\ast}c_{l}^{\ast}\max\left\{\frac{X}{M},\frac{1}{||\alpha ml||}\right\}.$
(Here and below, we treat the cases $l>0$ and $l<0$ in a similar way.) Writing $ml=k$, using the bound
$\alpha_{k}:=\sum_{\begin{subarray}{c}m\sim M\\ l\sim U\\ ml=k\end{subarray}}a_{m}^{\ast}c_{l}^{\ast}\ll X^{\varepsilon}$ (5.5)
and applying (5.2) with $K:=MU$ and $N:=X/M$, it follows that
$\mathcal{S}_{1}^{\ast}\ll\left(\frac{UX}{q}+MU+q\right)X^{\varepsilon}.$
Recalling
$U\leq L\ll X^{\theta+\eta}\quad\mbox{and}\quad X^{2\theta+10\eta}\leq q\leq X^{1-\theta-10\eta},$ (5.6)
the condition on $\theta$ in (5.4) and $M\leq X^{15/22}$ from (4.7), we deduce that
$\mathcal{S}_{1}^{\ast}\ll\left(X^{1-\theta-10\eta}+X^{15/22+\theta}\right)X^{\varepsilon+\eta}\ll X^{1-2\eta}$
if $\varepsilon$ and $\eta$ are suitably small, as desired.
To estimate $\mathcal{T}_{1}^{\ast}$, we define and bound $\alpha_{k}$ as in (5.5), set $K:=MU$ and $\beta_{n}:=b_{n}^{\ast}$ and recall $N\asymp X/M$ from (4.8), thus obtaining
$\mathcal{T}_{1}^{\ast}\ll(UX)^{1/2}\left(\frac{UX}{q}+MU+\frac{X}{M}+q\right)^{1/2}X^{\varepsilon}.$
Recalling (5.4), (5.6) and $X^{7/22}\leq M\leq X^{8/22}$ from (4.8), we deduce that
$\mathcal{T}_{1}^{\ast}\ll X^{(1+\theta)/2}\left(X^{1-\theta-10\eta}+X^{8/22+\theta}+X^{15/22}\right)^{1/2}X^{\varepsilon+\eta}\ll X^{1-2\eta}$
if $\varepsilon$ and $\eta$ are suitably small, as desired. ∎
6. Estimations of $\mathcal{S}_{2}^{\ast}$ and $\mathcal{T}_{2}^{\ast}$
Our estimations of $\mathcal{S}_{2}^{\ast}$ and $\mathcal{T}_{2}^{\ast}$ follow closely the treatments of similar type I and II sums in [7]. A crucial tool in this connection is the following standard estimate for exponential sums.
Lemma 6.1 (van der Corput).
Let $a$ and $b$ be integers such that $a<b$. Suppose that $f:[a,b]\rightarrow\mathbb{R}$ is a twice continuously differentiable function satisfying
$f^{\prime\prime}(t)\asymp\Lambda\quad\text{for all }t\in[a,b],$
where $\Lambda>0$. Then
$\left|\sum_{a<n\leq b}e(f(n))\right|\ll(b-a)\Lambda^{1/2}+\Lambda^{-1/2}.$
Proof.
This is [3, Theorem 2.2]. ∎
We first estimate the type II sum $\mathcal{T}_{2}^{\ast}$.
Lemma 6.2.
There is $\eta>0$ such that $\mathcal{T}_{2}^{\ast}\ll X^{1-2\eta}$.
Proof.
Estimating $\mathcal{T}_{2}^{\ast}$ in a similar way as the term $L$ in [7, section 4], we have
$\left|\mathcal{T}_{2}^{\ast}\right|^{2}\ll$ $\left(X^{5/2-\gamma}+X^{3-2\gamma}+X^{3-\gamma}N^{-1}+X^{(10-5\gamma)/3}N^{1/3}+X^{(8-4\gamma)/3}N^{2/3}+\right.$
$\left.X^{(10-8\gamma)/3}N^{4/3}\right)X^{2(\varepsilon+\eta)}$
$\ll$ $\left(X^{5/2-\gamma}+X^{3-2\gamma}+X^{2-\gamma}M+X^{(11-5\gamma)/3}M^{-1/3}+X^{(10-4\gamma)/3}M^{-2/3}+\right.$
$\left.X^{(14-8\gamma)/3}M^{-4/3}\right)X^{2(\varepsilon+\eta)},$
where the first line is found in [7, page 257] (with $X$ replaced by $N$, and $N$ replaced by $Y$), and for the second line we have used that $MN\asymp X$. Thus $\mathcal{T}_{2}^{\ast}\ll X^{1-2\eta}$ if
$\gamma>\frac{1}{2}+2\varepsilon+6\eta$ (6.1)
and
$X^{5-5\gamma+6\varepsilon+18\eta}\ll M\ll X^{\gamma-2\varepsilon-6\eta}.$ (6.2)
Interchanging the roles of $m$ and $n$, we also have $\mathcal{T}_{2}^{\ast}\ll X^{1-2\eta}$ under the conditions (6.1) and
$X^{5-5\gamma+6\varepsilon+18\eta}\ll N\asymp\frac{X}{M}\ll X^{\gamma-2\varepsilon-6\eta},$
i.e.
$X^{1-\gamma+2\varepsilon+6\eta}\ll M\ll X^{5\gamma-4-6\varepsilon-18\eta}.$ (6.3)
Recalling $X^{7/22}\leq M\leq X^{8/22}$ from (4.8) and $\gamma=1/c>8/9$ from Theorem 1.2, and noting that $41-\gamma<7/22<8/22<5\gamma-$ 4 if $\gamma>8/9$, the desired bound $\mathcal{T}_{2}^{\ast}\ll X^{1-2\eta}$ now follows upon taking $\varepsilon$ and $\eta$ sufficiently small. ∎
We point out that we have not used (6.2) but only (6.3) to establish Lemma 6.2. However, if $b_{n}^{\ast}=1$, we immediately deduce the following result on $\mathcal{S}_{2}^{\ast}$ under the condition (6.2).
Lemma 6.3.
Suppose that $\varepsilon$ and $\eta$ are small enough and (6.2) holds. Then $\mathcal{S}_{2}^{\ast}\ll X^{1-2\eta}$.
This will be useful if $M$ is large. Next, we prove the following for $M$ in a medium range.
Lemma 6.4.
Suppose that $\varepsilon$ and $\eta$ are small enough and
$X^{2-2\gamma+2\varepsilon+6\eta}\ll M\ll X^{\min\{2/3,4\gamma-3-4\varepsilon-12\eta\}}.$
Then $\mathcal{S}_{2}^{\ast}\ll X^{1-2\eta}$.
Proof.
Following the estimation of the term $K$ in [7, section 5] and taking the exponent pair $(p,q)=(1/2,1/2)$ in [7, Lemma 7], we have
$\mathcal{S}_{2}^{\ast}\ll X^{-\gamma}\left(X^{2}N^{-1/4}+X^{3/2}N^{1/2}+X^{7/4}N^{1/8}\right)X^{\varepsilon+\eta},$
provided that $N\geq X^{1/3}$. (Again, $X,N$ replace the variables $N,Y$ in [7], respectively.) Using $MN\asymp X$, it follows that
$\mathcal{S}_{2}^{\ast}\ll\left(X^{7/4-\gamma}M^{1/4}+X^{2-\gamma}M^{-1/2}+X^{15/8-\gamma}M^{-1/8}\right)X^{\varepsilon+\eta},$
provided that $M\leq X^{2/3}$. This implies the result in Lemma 6.4. ∎
Finally, for small $M$, we prove the following by a direct appeal to Lemma 6.1.
Lemma 6.5.
Suppose that $\eta$ is small enough and $M\ll X^{\gamma-1/2-4\eta}$. Then $\mathcal{S}_{2}^{\ast}\ll X^{1-2\eta}$.
Proof.
Applying Lemma 6.1 with $f(n):=h(mn)^{\gamma}$ to bound the smooth sum over $n$, and summing over $m$ and $h$ trivially, we obtain
$\mathcal{S}_{2}^{\ast}\ll V^{3/2}X^{\gamma/2}M+V^{1/2}X^{1-\gamma/2}\ll X^{3/2-\gamma+2\eta}M,$
where we have used $V\leq H\ll X^{1-\gamma+\eta}$. This implies the result in Lemma 6.5. ∎
Since
$(0,\gamma-1/2)\cup(2-2\gamma,\min\{2/3,4\gamma-3\})\cup(5-5\gamma,\gamma)=(0,\gamma)\supset(0,15/22)$
if $\gamma>8/9$, Lemmas 6.3, 6.4 and 6.5 cover all relevant ranges if $\varepsilon$ and $\eta$ are small enough, and we thus have the following.
Lemma 6.6.
There is $\eta>0$ such that $\mathcal{S}_{2}^{\ast}\ll X^{1-2\eta}$.
Here we point out that the condition $\gamma>8/9$ above comes from the inequality $35-5\gamma<4\gamma-$ 3.
7. Estimations of $\mathcal{S}_{3}^{\ast}$ and $\mathcal{T}_{3}^{\ast}$
Our estimations of $\mathcal{S}_{3}^{\ast}$ and $\mathcal{T}_{3}^{\ast}$ follow closely the treatments of similar type I and II sums in [1]. We first estimate the type II sum $\mathcal{T}_{3}^{\ast}$.
Lemma 7.1.
There is $\eta>0$ such that $\mathcal{T}_{3}^{\ast}\ll X^{1-2\eta}$.
Proof.
Estimating the triple sum over $m,n,h$ on the right-hand side of (4.6) similarly as in [1, inequality (4.4)], and summing over $l$ trivially, we obtain
$\begin{split}\mathcal{T}_{3}^{\ast}\ll U&\left(M^{1/2}Q^{1/2}V^{1/2}X^{1/2}+V^{5/4}X^{1+\gamma/4}M^{-1/2}Q^{-1/4}+Q^{1/2}V^{1/2}X^{1-\gamma/2}+\right.\\ &\left.Q^{1/4}V^{3/4}X^{1-\gamma/4}\right)X^{\varepsilon}\end{split}$ (7.1)
if $Q\gg 1$. We choose $Q$ in such a way that the first term satisfies $UM^{1/2}Q^{1/2}V^{1/2}X^{1/2+\varepsilon}\ll X^{1-2\eta}$, i.e.
$Q:=\frac{X^{1-2\varepsilon-4\eta}}{U^{2}MV}.$ (7.2)
Using $U\leq L\ll X^{\theta+\eta}$ and $V\leq H\ll X^{1-\gamma+\eta}$, the condition $Q\gg 1$ holds if
$M\ll X^{\gamma-2\theta-2\varepsilon-7\eta}.$ (7.3)
Plugging (7.2) into (7.1), and using $U\ll X^{\theta+\eta}$ and $V\ll X^{1-\gamma+\eta}$ again, we get
$\mathcal{T}_{3}^{\ast}\ll$ $X^{1-2\eta}+\left(U^{3/2}V^{3/2}X^{3/4+\gamma/4}M^{-1/4}+X^{3/2-\gamma/2}M^{-1/2}+U^{1/2}V^{1/2}X^{5/4-\gamma/4}M^{-1/4}\right)X^{2\epsilon+\eta}$
$\ll$ $X^{1-2\eta}+\left(X^{9/4+3\theta/2-5\gamma/4}M^{-1/4}+X^{3/2-\gamma/2}M^{-1/2}+X^{7/4+\theta/2-3\gamma/4}M^{-1/4}\right)X^{2\varepsilon+3\eta}.$
The last line is $\ll X^{1-2\eta}$ if
$M\gg X^{5-5\gamma+6\theta+8\varepsilon+20\eta}.$ (7.4)
Combining (7.3) and (7.4), the desired bound $\mathcal{T}_{3}^{\ast}\ll X^{1-2\eta}$ holds provided that
$X^{5-5\gamma+6\theta+8\varepsilon+20\eta}\ll M\ll X^{\gamma-2\theta-2\varepsilon-7\eta}.$ (7.5)
Interchanging the roles of $m$ and $n$, we also have $\mathcal{T}_{3}^{\ast}\ll X^{1-2\eta}$ when
$X^{5-5\gamma+6\theta+8\varepsilon+20\eta}\ll N\asymp\frac{X}{M}\ll X^{\gamma-2\theta-2\varepsilon-7\eta},$
i.e.
$X^{1-\gamma+2\theta+2\varepsilon+7\eta}\ll M\ll X^{5\gamma-4-6\theta-8\varepsilon-20\eta}.$ (7.6)
Recalling our condition $X^{7/22}\leq M\leq X^{8/22}$ from (4.8) and observing that
$1-\gamma+2\theta<\frac{7}{22}<\frac{8}{22}<5\gamma-4-6\theta$
under the conditions in (5.4), the desired bound $\mathcal{T}_{3}^{\ast}\ll X^{1-2\eta}$ now follows upon taking $\varepsilon$ and $\eta$ sufficiently small. ∎
We point out that we have not used (7.5) but only (7.6) to establish Lemma 7.1. However, if $b_{n}^{\ast}=1$, we immediately deduce the following result on $\mathcal{S}_{3}^{\ast}$ under the condition (7.5).
Lemma 7.2.
Suppose that $\varepsilon$ and $\eta$ are small enough and (7.5) holds. Then $\mathcal{S}_{3}^{\ast}\ll X^{1-2\eta}$.
This will be useful if $M$ is large. Next, we prove the following for $M$ in a medium range.
Lemma 7.3.
Suppose that $\varepsilon$ and $\eta$ are small enough and
$X^{7-8\gamma+8\theta+8\varepsilon+40\eta}\ll M\ll X^{4\gamma-3-4\theta-4\varepsilon-20\eta}.$
Then $\mathcal{S}_{3}^{\ast}\ll X^{1-2\eta}$.
Proof.
Estimating the triple sum over $m,n,h$ on the right-hand side of (4.3) similarly as in [1, last inequality on page 60], and summing over $l$ trivially, we obtain
$\mathcal{S}_{3}^{\ast}\ll U\left(V^{5/4}M^{1/4}X^{1/2+\gamma/4}+V^{7/8}M^{-1/8}X^{1-\gamma/8}\right)X^{\varepsilon}.$
Using $U\ll X^{\theta+\eta}$ and $V\ll X^{1-\gamma+\eta}$, it follows that
$\mathcal{S}_{3}^{\ast}\ll\left(X^{7/4+\theta-\gamma}M^{1/4}+X^{15/8+\theta-\gamma}M^{-1/8}\right)X^{\varepsilon+3\eta}.$
This implies the result in Lemma 7.3. ∎
Finally, for small $M$, we prove the following by a direct appeal to Lemma 6.1.
Lemma 7.4.
Suppose that $\eta$ is small enough and $M\ll X^{\gamma-1/2-\theta-5\eta}$. Then $\mathcal{S}_{3}^{\ast}\ll X^{1-2\eta}$.
Proof.
Applying Lemma 6.1 with $f(n):=h(mn)^{\gamma}$ to bound the smooth sum over $n$, and summing over $m$, $l$ and $h$ trivially, we obtain
$\mathcal{S}_{3}^{\ast}\ll U\left(V^{3/2}X^{\gamma/2}M+V^{1/2}X^{1-\gamma/2}\right)\ll X^{3/2-\gamma+\theta+3\eta}M,$
where we have used $U\ll X^{\theta+\eta}$ and $V\ll X^{1-\gamma+\eta}$. This implies the result in Lemma 7.4. ∎
Since
$(0,\gamma-1/2-\theta)\cup(7-8\gamma+8\theta,4\gamma-3-4\theta)\cup(5-5\gamma+6\theta,\gamma-2\theta)=(0,\gamma-2\theta)\supset(0,15/22)$
under the conditions in (5.4), Lemmas 7.2, 7.3 and 7.4 cover all relevant ranges if $\varepsilon$ and $\eta$ are small enough, and we thus have the following, completing the proof of Theorem 1.2.
Lemma 7.5.
There is $\eta>0$ such that $\mathcal{S}_{3}^{\ast}\ll X^{1-2\eta}$.
Here we point out that the condition $\theta<(9\gamma-8)/10$ in (5.4) comes from the inequality $5-5\gamma+6\theta<4\gamma-3-4\theta$. We also note that $\mathcal{S}_{2}^{\ast}$ and $\mathcal{T}_{2}^{\ast}$ could have been bounded in a similar way as $\mathcal{S}_{3}^{\ast}$ and $\mathcal{T}_{3}^{\ast}$ above since the method of Balog and Friedlander also applies to the case when $l=0$. However, we decided to keep our separate treatment of $\mathcal{S}_{2}^{\ast}$ and $\mathcal{T}_{2}^{\ast}$ because the condition $\gamma>8/9$ emerges most naturally from it, and these terms are independent of $\theta$.